PLEASE Check my Work !
Among the countries , Nigeria had the highest population growth of 62.4% . Nigeria's pollution is 97,552 thousands in 1990 . Find the population of Nigeria in 2010 .Round to the nearest thousands .
62.4/100 = 0.624 =1.624
1.624x = 97,552
x = 97,552 / 1.624
x = 60,068.96
x= 60,069 nearest thousands

Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)

Nigeria had the highest population growth of 62.4%
Nigeria's pollution is 97,552 thousands in 1990.
Find the population of Nigeria in 2010 (x)
 
Hi,
Let x be the population of Nigeria in 2010.
Had the right idea on the 1.624. Good Job.
However it was the 997,552 that was doing the increasing.
So that would be:
x = 1.624(97,552) = 158,424 the population of Nigeria in 2010
Wish You the Best in your Studies.

Answer by ikleyn(53763)

.
Among the countries , Nigeria had the highest population growth of 62.4% . Nigeria's pollution is 97,552 thousands in 1990.
Find the population of Nigeria in 2010. Round to the nearest thousands.
62.4/100 = 0.624 =1.624
1.624x = 97,552
x = 97,552 / 1.624
x = 60,068.96
x= 60,069 nearest thousands.
~~~~~~~~~~~~~~~~~


            Your solution is  ABSOLUTELY  WRONG.

            Think a minute: the population   G R O W S,   but your final number is  LESS  than starting number (!)
            It is  NONSENSE (!)

            Below see the  CORRECT  SOLUTION.     It is in one line.


The populatin of Nigeria in 2010, according to your input, is (or was)  97,552 + 0.624*97,552 = 158,454 thousands.

Solved, answered and explained.


About the population of  Nigeria,  see this  Wikipedia article

https://www.google.com/search?q=population+of+nigeria&rlz=1C1CHBF_enUS910US910&oq=population+of+Nigeria&aqs=chrome.0.0i433j0l3j0i395l6.8559j1j15&sourceid=chrome&ie=UTF-8


============

On percentage,  see the lessons
    - Percentage problems
    - Percentage word problems (Type 1 problems, Finding the Part)
    - Percentage word problems (Type 2 problems, Finding the Rate)
    - Percentage word problems (Type 3 problems, Finding the Base)
    - Simple percentage problems
    - More complicated percentage problems
    - Advanced problems on percentage
    - Percentage problems on chains of discounts
    - Problems on percentage that lead to unexpected results
    - Simple interest percentage problems
    - Compounded interest percentage problems
    - OVERVIEW of lessons on percentage problems
in this site.

They are written specially for you.

Read them and become an expert in this area.

Start from the very first lesson of the list.

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.